Hey,
Please help me solve this problem:
A square mat used for athletic exercises has a uniform red border on all four sides. The rest of the mat is blue. The width of the blue square is three-fourths the width of the entire square. If the area colored red is 28 m^2, determine the length of each side of the mat.
Thanks!

Answer by scott8148(6628)

let x=length ... (whole mat)-(blue area)=(red area) ... x^2-((3/4)x)^2=28

x^2-(9/16)x^2=28 ... (7/16)x^2=28